Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used iRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the predetermined overhead rate for the year. Predetermined ovRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the amount of underapplied or overapplied overhead for the yeaRequired 1 Required 2 Required 3 Required 4 Required 5 Prepare a schedule of cost of goods manufactured for the year. AssumeRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the unadjusted cost of goods sold for the year. Do not includeRequired 1 Required 2 Required 3 Required 4 Required 5 Assume that the $39,000 ending balance in Work in Process includes $8,
Answer #1
Required 1 :
Predetermined overhead rate = Estimated manufacturing overhead / Estimated direct materials costs = 111800 / 86000 130%
Required 2 :
Beginning raw materials 23000
(+) Purchase of raw materials 133000
Total raw materials available 156000
(-) Ending raw materials 13000
Raw materials used 143000
Overhead applied = Raw materials used * Predetermined overhead rate = 143000 * 130% 185900
Indirect labor 108600
Property taxes 8800
Depreciation of equipment 18000
Maintenance 14000
Insurance 10900
Rent, building 30000
Actual overhead costs 190300
As the overhead applied are less than the actual overhead costs, the overheads are underapplied
Underapplied overhead = Actual overhead costs - Overhead applied = 190300 - 185900 4400
Underapplied Overhead 4400
Required 3 :
Gitano Products
Schedule of cost of goods manufactured
Direct materials :
Beginning raw materials 23000
Purchase of raw materials 133000
   Total raw materials available 156000
Ending raw materials 13000
Raw materials used in production 143000
Direct labor cost 82000
Manufacturing overhead applied 185900
Total manufacturing costs 410900
Beginning work in process 50000
Total cost of work in process 460900
Ending work in process 39000
Cost of goods manufactured 421900
Required 4 :
Cost of goods manufactured 421900
(+) Beginning finished goods 72000
(-) Ending finished goods 58000
Unadjusted cost of goods sold 435900
Required 5 :
Direct materials 8400
Direct labor cost ( 39000 - 8400 - 10920 ) 19680
Manufacturing overhead ( 8400 * 130% ) 10920
Work in process inventory 39000